# WorldGUI: An Interactive Benchmark for Desktop GUI Automation from Any Starting Point

### v3

@@ D VSCode, File Explorer Word, Adobe Acrobat
Creating Augmented Tasks. In our study, to stimulate dynamic testing processing in real GUI interactions, we propose to design GUI tasks with various initial tasks. Specifically, we propose pre-actions before executing the task. The pre-actions primarily serve two purposes: 1) Simulating Intermediate Task States: Pre-actions can complete specific steps of a task, creating a starting point from an intermediate state. This approach addresses scenarios where users may seek AI assistance because they are unable to complete a task. For example, if the task involves opening a dropdown menu, the pre-action may pre-open the menu. If the agent fails to recognize this precondition and follows its plan to click the menu again, it might inadvertently close the menu, causing task failure.
Introducing Diverse Initial Context States: Pre-actions can also introduce variations in the initial state, such as opening random tabs or settings. This ensures that the starting state is unconventional, challenging the agent to adapt by modifying its plan or adding new steps. We illustrate one example in Figure 8. Here, the meta task and augmented task, have the same user query and instructional video and it will ideally have the same final state. We additionally provide more examples about augmenting the meta task in Figure 9.
@@
Appendix C Data Statistics
Augmentation tasks type analysis. As we summarize, the real GUI scenarios include: (1) The software interface is not in its default state. (2) The human-computer interactions may start from the intermediate state of a specific task. We propose to create the augmented tasks for each meta task to stimulate authentic GUI interactions. Our augmentations lie in two main groups: (1) stimulating the intermediate states and (2) introducing diverse initial states. We divide the two groups into three exact types: Add-step, Trim-step, and Adjust-step. For Add-step, it represents various unrelated state augmentations to stimulate the scenario that we may start the agent-computer interactions in another unrelated task or interfaces, the agent should replan the task to add necessary steps. For Trim-step, it represents that we finish several steps of a long task and make the task in an intermediate state. For Adjust-step, it is usually a small modification of the existing state, such as changing the interface by clicking another Tab or clicking a button to open an unrelated dropdown menu. Most of the time, it would not require new steps to return to the target task progress. This augmentation may mislead the agent in state understanding, making them jump or miss the key steps. As shown in Figure 11, the manually created augmentations mainly belong to the add-step. Adjust-step could be the second-largest application, except for the File Explorer. Due to the low complexity of the interfaces of File Explorer, we cannot create many augmentations for adjust-step.
Task difficulty analysis. Figure 12 shows the distribution of the task difficulty across desktop applications. We annotated the task difficulty level based the subjective software usage experience. The results indicate that the tasks in Adobe Acrobat and VLC player are more challenging. The tasks in Excel, PowerPoint, and Word are more at the medium and simple levels. By considering the Success Rate and task length on these tasks, one can know that the tasks are easy for humans, but hard for current GUI agents. According to VSCode and File Explorer, and YouTube applications, the tasks are easier than in other applications. Overall, the task difficulty of created data is diverse across different applications, and there is still a need for stronger agents focusing on handling desktop-oriented GUI tasks.
The task details about the user query and the pre-actions are included in the metadata JSON file in the supplementary materials. The project file, instruction video, and augmentation files can be found in the provided data link.
Figure 10: Distribution of Software taxonomy and the distribution of task length.
Figure 11: The distribution of different augmentation types.
Figure 12: The distribution of different task difficulty.
Table 10: Task statistics of WorldGUI.
Category Software Meta Count Aug Count Meta-Task Length (Avg.)
Office Adobe Acrobat 11 55 7.1
Office Excel 12 58 4.1
Office PowerPoint 11 53 5.1
Office Word 11 52 5.1
Coding VSCode 11 45 4.4
Windows Usage File Explorer 10 34 7.1
Windows Usage Settings 14 55 5.1
Web Usage Web Browser 10 49 7.8
Web Usage YouTube 11 50 4.8
Media VLC Player 10 49 10.8
Total (Average) 111 500 6.0
Appendix D Detailed Experimental Results
Table 11 shows the detailed results of WorldGUI-Agent across individual applications in WindowsAgentArena [4] benchmark. The results of this related Windows-centric interactive GUI benchmark indicate that current the desktop GUI tasks are more challenging than web tasks. As we complete 11 out of 17 tasks in Web Browsing, a similar phenomenon is also discovered in Table 4.
Table 11: Detailed experimental results of WorldGUI-Agent across individual applications in WindowsAgentArena [4].
@@ Appendix D Detailed Experimental Results
Appendix E Computational Costs Discussion
@@
Take a Windows Setting task as an example, we provide detailed time costs across different modules tested on a Laptop with AMD Ryzen 7 4800H CPU in Table 13. Task length: 6 (Generated by Planner+Planner-Critic). Total steps: 18, total time: 335.84s, total agent execution time: 412.37s. Since desktop GUI is in its early stages, computational costs are currently unavoidable. Even OpenAI’s Deep Research takes over 10 minutes in daily use. According to the post from OpenAI’s Operator, achieving 38.1% on OS-World requires over 100 steps, which is costly as well. In summary, it is still a tradeoff between performance and time costs in GUI automation.
Table 13: Time costs of each module.
@@
0 Planner+Planner-Critic 12.53
1 Step-Check+Actor 21.23
1 Actor-Critic 13.07
2 Step-Check+Actor 17.81
2 Actor-Critic 11.86
3 Step-Check+Actor 18.43
3 Actor-Critic 11.72
4 Step-Check+Actor 22.49
4 Actor-Critic 25.31
4 Actor-Critic 10.82
5 Step-Check+Actor 17.01
5 Actor-Critic 22.36
5 Actor-Critic 26.85
5 Actor-Critic 18.99
6 Step-Check+Actor 21.79
6 Actor-Critic 23.37
6 Actor-Critic 21.12
6 Actor-Critic 19.08
Appendix F Examples of Augmentations
In this section, we present several augmentation examples in Figures 13, 14, 15, 16, 17, 18, 19. It is noted that our augmentations are not only making the first step changing but also require the agent add new step in its second step. For instance, in Figure 13, our augmentation is about click on Data tab in the ribbon, in the default software state, the Merge & Center button exhibit in the Home tab, there is no need to click on Home tab, after our augmentations, the agent should add a new task “Click on Home Tab” before it click on the Merge & Center button. Similarly, in Figure 18, the text editing buttons are under the Home Tab, if we augment the initial state with other Tab like Animation Tab, after the first step “Select the text ’US SUBMARINE DAY’ ”, the agent should add a new step like “Click on Home Tab” back to the default state for task execution. Except for adding new steps, we also present an example about adjust the step in Figure 14, as the target is about merging cells A1 to K1, we augment the initial state by selecting A2 to K2. Such a slight difference may mislead the agent to perceive such a minor difference, and the agent may jump the first step about selecting the correct cells lead to finally unsucess. In Figure 15 and Figure 19, we show two examples of introducing pop-up window in the initial state which require the agents accurately identify the pop-up windows and correctly close it by replanning the task based on the visual screenshot not only strictly planning based on inherited knowledge or the instructinal videos. In Figure 16, we show an example of changing the interface by clicking the Data tab to hide the Merge & Center button under the Home tab. In Figure 17, we complete the first step about selecting A1 to K1, which requires the agent to jump this step to reduce the time costs.
@@ Appendix F Examples of Augmentations
Appendix G WorldGUI-Agent Reasoning Loop Algorithm
@@
  Initial current subtask $S_{t=0}=S_{1}^{1}$, where $S_{1}^{0}$ is the $1$-th subtask in the $1$-th milestone of $p$.
  Initial $s_{0}=\textless Continue\textgreater$ >
@@
     if $s_{t}=\textless Next\textgreater$ > then
@@
     Set $C_{t}=$ None; $t=t+1$; Set state $s_{t}=\textless Critic\textgreater$ > (For each subtask, the first step is finished, then execute the actor-critic process)
@@
     if $s_{t}=\textless Next\textgreater$ > then
@@
     while $s_{t}=\textless Critic\textgreater$ > and $z<$ max critique trials do
@@
        if $s_{t}=\textless Next\textgreater$ > then
@@
Appendix H Qualitative Results
(1) In Figure 20, we present a successful prediction example, demonstrating that the WorldGUI can effectively plan each step for a task, accurately perceive specific elements in the GUI, and convert them into the correct action code. Additionally, we display the parsed GUI elements, which can accurately identify most content, including small icons and dense text elements.
@@ Appendix H Qualitative Results
(2) We provide the visualization results of using Planner-Critic, Step-Check, and Actor-Critic in Figure 22, Figure 23, and Figure 24. These qualitative results demonstrate the effectiveness of these critical modules in GUI automation.
@@
(3) We also highlight some common errors encountered. 1) The model has the difficulty of obtaining the desired information when we augment the task by invoking the dropdown menu of the Settings application. As shown in the left of Figure 25, when we click on the ’System’ button on the left, it is challenging for our model to extract the button’s position as it is hidden. Such cases require the model to have a higher level of ability to delete the content in the input box or click on the blank area. 2) As shown in the right of Figure 25, the model has difficulty dragging a bar to achieve the desired value. 3) The model struggles with the visual choice when there is no text information in the screenshot, as shown on the left of Figure 26. The subtask aims to select the center button, but the current model makes it hard to detect the center choice only from the screenshot. 4) The model cannot successfully locate the position of the input box, as the GUI parser will easily locate the text location ’Replace with’, it always outputs the action like clicking on the ’Replace with’, which will destroy the whole task success.

### v4

@@ D VSCode, File Explorer Word, Adobe Acrobat
D.2 Creating Augmented Tasks
In our study, to simulate dynamic testing processing in real GUI interactions, we propose to design GUI tasks with various initial tasks. Specifically, we propose pre-actions before executing the task. The pre-actions primarily serve two purposes: 1) Simulating Intermediate Task States: Pre-actions can complete specific steps of a task, creating a starting point from an intermediate state. This approach addresses scenarios where users may seek AI assistance because they are unable to complete a task. For example, if the task involves opening a dropdown menu, the pre-action may pre-open the menu. If the agent fails to recognize this precondition and follows its plan to click the menu again, it might inadvertently close the menu, causing task failure.
D.3 Introducing Diverse Initial Context States
Pre-actions can also introduce variations in the initial state, such as opening random tabs or settings. This ensures that the starting state is unconventional, challenging the agent to adapt by modifying its plan or adding new steps. We illustrate one example in Figure 5. Here, the meta task and augmented task, have the same user query and instructional video and it will ideally have the same final state. We additionally provide more examples about augmenting the meta task in Figure 11.
@@ D.3 Introducing Diverse Initial Context States
Appendix E Detailed Experimental Results
Table 14 shows the detailed results of WorldGUI-Agent across individual applications in WindowsAgentArena (Bonatti et al., 2024) benchmark. The results of this related Windows-centric interactive GUI benchmark indicate that current the desktop GUI tasks are more challenging than web tasks. As we complete 11 out of 17 tasks in Web Browsing, a similar phenomenon is also discovered in Table 4.
Table 14: Detailed experimental results of WorldGUI-Agent across individual applications in WindowsAgentArena (Bonatti et al., 2024).
@@ Appendix E Detailed Experimental Results
Appendix F Computational Costs Discussion
@@
Take a Windows Setting task as an example, we provide detailed time costs across different modules tested on a desktop PC with AMD Ryzen 7 5800H CPU. Task length: 6 (generated by Planner+Planner-Critic). To facilitate a fair comparison, we additionally selected two of the latest SOTA agents, Agent-S (2024-10-08) and Agent-S2 (2025-04-01), and measured their run times on the same successful task under identical hardware and the same base MLLM (Claude-Sonnet-4). The results are shown in Tables 16, 17, 18. To summarize, our WorldGUI-Agent shows a competitive running time of 129.55s, as compared with Agent-S (131.98s) and Agent-S2 (108.64s). The main computational costs of our designed modules are largely affected by the underlying large multimodal model, leaving room for acceleration optimization.
Since desktop GUI automation is still in its early stages, such computational costs are currently unavoidable. For reference, even OpenAI’s Deep Research reportedly takes over 10 minutes in daily usage. According to OpenAI Operator’s report, achieving 38.1% on OS-World requires over 100 steps, which is similarly costly. In summary, there remains a clear tradeoff between performance and time costs in GUI automation, and this challenge is shared across the community.
Table 16: Running time with WorldGUI-Agent (ours).
@@
0 Planner 4.48
0 Planner-Critic 11.47
1 Parser 2.03
1 Step-Check 4.95
1 Actor 7.47
1 Parser 2.07
1 Actor-Critic 7.38
2 Parser 2.03
2 Step-Check 6.71
2 Actor 6.05
2 Parser 2.07
2 Actor-Critic 7.97
3 Parser 2.12
3 Step-Check 6.35
3 Actor 6.71
3 Parser 2.22
3 Actor-Critic 10.06
4 Parser 2.01
4 Step-Check 6.19
4 Actor 8.54
4 Parser 2.40
4 Actor-Critic 9.55
5 Parser 2.22
5 Step-Check 6.50
Total - 129.55
Average (per action) - 22.72
Table 17: Running time with Agent-S.
Step Executed Modules Time (seconds)
0 Manager 2.02
1 Manager 10.11
2 Worker 16.05
3 Worker 12.38
4 Worker 17.30
5 Worker 12.26
6 Worker 15.17
7 Worker 15.46
8 Worker 11.83
9 Worker 19.40
Total - 131.98
Average (per worker) - 14.98
Table 18: Running time with Agent-S2.
Step Executed Modules Time (seconds)
0 Manager 11.87
1 Worker 7.27
2 Worker 13.21
3 Worker 13.59
4 Worker 14.04
5 Manager 10.61
6 Worker 9.87
7 Manager 7.53
8 Worker 20.65
Total - 108.64
Average (per worker) - 13.22
Appendix G Examples of Augmentations
In this section, we present several augmentation examples in Figures 12, 13, 14, 15, 16, 17, 18. It is noted that our augmentations are not only making the first step changing but also require the agent add new step in its second step. For instance, in Figure 12, our augmentation is about click on Data tab in the ribbon, in the default software state, the Merge & Center button exhibit in the Home tab, there is no need to click on Home tab, after our augmentations, the agent should add a new task “Click on Home Tab” before it click on the Merge & Center button. Similarly, in Figure 17, the text editing buttons are under the Home Tab, if we augment the initial state with other Tab like Animation Tab, after the first step “Select the text ’US SUBMARINE DAY’ ”, the agent should add a new step like “Click on Home Tab” back to the default state for task execution. Except for adding new steps, we also present an example about adjust the step in Figure 13, as the target is about merging cells A1 to K1, we augment the initial state by selecting A2 to K2. Such a slight difference may mislead the agent to perceive such a minor difference, and the agent may jump the first step about selecting the correct cells lead to finally unsucess. In Figure 14 and Figure 18, we show two examples of introducing pop-up window in the initial state which require the agents accurately identify the pop-up windows and correctly close it by replanning the task based on the visual screenshot not only strictly planning based on inherited knowledge or the instructinal videos. In Figure 15, we show an example of changing the interface by clicking the Data tab to hide the Merge & Center button under the Home tab. In Figure 16, we complete the first step about selecting A1 to K1, which requires the agent to jump this step to reduce the time costs.
@@ Appendix G Examples of Augmentations
Appendix H WorldGUI-Agent Reasoning Loop Algorithm
@@
 Initial current subtask $S_{t=0}=S_{1}^{1}$, where $S_{1}^{0}$ is the $1$-th subtask in the $1$-th milestone of pp.
 Initial $s_{0}=\textless Continue\textgreater$
@@
  if $s_{t}=\textless Next\textgreater$ then
@@
  Set $C_{t}=$ None; $t=t+1$; Set state $s_{t}=\textless Critic\textgreater$ (For each subtask, the first step is finished, then execute the actor-critic process)
@@
  if $s_{t}=\textless Next\textgreater$ then
@@
  while $s_{t}=\textless Critic\textgreater$ and $z<$ max critique trials do
@@
   if $s_{t}=\textless Next\textgreater$ then
@@
Appendix I Qualitative Results
(1) In Figure 19, we present a successful prediction example, demonstrating that the WorldGUI can effectively plan each step for a task, accurately perceive specific elements in the GUI, and convert them into the correct action code. Additionally, we display the parsed GUI elements, which can accurately identify most content, including small icons and dense text elements. (2) We provide the visualization results of using Planner-Critic, Step-Check, and Actor-Critic in Figure 21, Figure 22, and Figure 23. These qualitative results demonstrate the effectiveness of these critical modules in GUI automation. (3) We also highlight some common errors encountered. 1) The model has difficulty obtaining the desired information when we augment the task by invoking the dropdown menu of the Settings application. As shown on the left of Figure 24, when we click on the ’System’ button on the left, it is challenging for our model to extract the button’s position as it is hidden. Such cases require the model to have a higher level of ability to delete the content in the input box or click on the blank area. 2) As shown in the right of Figure 24, the model has difficulty dragging a bar to achieve the desired value. 3) The model struggles with the visual choice when there is no text information in the screenshot, as shown on the left of Figure 25. The subtask aims to select the center button, but the current model makes it hard to detect the center choice only from the screenshot. 4) The model cannot successfully locate the position of the input box, as the GUI parser will easily locate the text location ’Replace with’, it always outputs the action like clicking on the ’Replace with’, which will destroy the whole task’s success.